Divide
4d exponent 9 m exponent 6 * 10d exponent 9 m exponent 6
__________________________________________________________
2d exponent 9 m exponent 9

Answer by jayanthihemkos@yahoo.com(87)

4d exponent 9 m exponent 6 * 10d exponent 9 m exponent 6/2d exponent 9 m exponent 9
solution:
40d^18m^12/2d^9m^9 ( 4*10 = 40 , d^9*d^9 = d^9+9 , m^6*m^6 = m^12)
20d^9 m^3 (20/2 = 20 , d^18/d^9 = d^18-9 = d^9, m^12/m^9 = m^12-9 =m^3)
Hence the solution